### v1

# MTIL: Encoding Full History with Mamba for Temporal Imitation Learning

Abstract
Standard imitation learning (IL) methods have achieved considerable success in robotics, yet often rely on the Markov assumption, limiting their applicability to tasks where historical context is crucial for disambiguating current observations. This limitation hinders performance in long-horizon sequential manipulation tasks where the correct action depends on past events not fully captured by the current state. To address this fundamental challenge, we introduce Mamba Temporal Imitation Learning (MTIL), a novel approach that leverages the recurrent state dynamics inherent in State Space Models (SSMs), specifically the Mamba architecture. MTIL encodes the entire trajectory history into a compressed hidden state, conditioning action predictions on this comprehensive temporal context alongside current multi-modal observations. Through extensive experiments on simulated benchmarks (ALOHA, Robomimic, LIBERO) and on challenging real-world sequential manipulation tasks that requiring long-term temporal reasoning, MTIL significantly outperforms state-of-the-art methods like ACT and Diffusion Policy. Our findings affirm the necessity of full temporal context for robust sequential decision-making and validate MTIL as a powerful approach that transcends the inherent limitations of Markovian imitation learning. Our implementation is available at MTIL
Keywords: Imitation Learning, Robot Manipulation, Sequence Modeling, State Space Models
1 Introduction
Imitation Learning (IL) has emerged as a powerful paradigm for teaching robots complex skills directly from expert demonstrations, bypassing the need for intricate reward engineering often required in reinforcement learning [1, 2, 3]. Behavioral Cloning (BC), the simplest form of IL, learns a direct mapping from observations to actions via supervised learning and has enabled robots to perform a variety of tasks [2, 3, 4, 1]. Recent advancements, particularly leveraging powerful sequence models and generative approaches, have led to state-of-the-art (SOTA) methods such as the Action Chunking Transformer (ACT) [5, 6] and Diffusion Policy [7, 8, 9, 10], which excel at learning visuomotor control policies for complex manipulation.
Despite these successes, a fundamental limitation persists in many current IL approaches: the reliance on the Markov assumption. These methods typically predict the action $a_{t}$ based solely on the current observation $o_{t}$ or a short, fixed-length history window $o_{t-k:t}$  [1, 11, 12, 13]. This assumption breaks down in tasks where the history beyond this limited window is necessary to resolve ambiguity in the current observation. Consider a sequential task requiring a robot to first place an object at location A, and subsequently move it to location B. At an intermediate configuration, the robot’s visual observation and proprioceptive state ($o_{t}$) might be identical regardless of whether it has successfully completed the sub-task at location A. A Markovian policy, lacking the memory of visiting A, cannot distinguish these fundamentally different historical contexts and may erroneously proceed directly to B, failing the task [1, 14, 15]. This temporal ambiguity signifies an underlying Partially Observable Markov Decision Process (POMDP), posing a critical challenge for standard IL methods in state-dependent tasks.
Since ambiguous tasks manifest as POMDPs and human demonstrations are inherently non-Markovian, effective imitation necessitates history-aware policies.While the critical role of historical context has been increasingly recognized  [12], effectively and efficiently encoding comprehensive trajectory histories into actionable representations remains a significant challenge. Addressing this, we introduce Mamba Temporal Imitation Learning (MTIL).This approach is specifically designed to incorporate the complete observational history into decision-making by leveraging the unique properties of State Space Models (SSMs), particularly the recently developed Mamba architecture [16, 17]. Mamba’s recurrent structure allows it to maintain a compressed hidden state $h_{t}$ that theoretically encapsulates information from the entire preceding observation sequence $H_{t}=(o_{1},...,o_{t})$. Instead of relying solely on $o_{t}$, MTIL learns a policy $\pi(a_{t}|h_{t},o_{t})$ that explicitly conditions the action prediction on this history-infused hidden state $h_{t}$ in conjunction with the current observation $o_{t}$, enabling differentiation between observationally similar states and the correct execution of complex sequential tasks. Our contributions are threefold:
  1. 1.
    A novel Mamba-based IL architecture, MTIL, is proposed, specifically designed for full trajectory history encoding via its recurrent state to overcome limitations inherent in Markovian sequential task learning.
  2. 2.
    The fundamental challenge of temporal ambiguity in IL is addressed by incorporating comprehensive historical context, offering a promising pathway towards learning and executing long-horizon, state-dependent manipulation sequences previously intractable for methods reliant on limited context.
  3. 3.
    Extensive empirical validation is provided across diverse benchmarks (simulation and real-world), demonstrating significant performance improvements over existing state-of-the-art methods (including ACT, Diffusion Policy), particularly on tasks demanding temporal reasoning and long-term memory.
2 Related Works
2.1 Markovian and Short-History Imitation Learning
A cornerstone of imitation learning, Behavioral Cloning (BC), typically learns a Markovian policy $\pi(a_{t}|o_{t})$ via supervised learning [2, 3, 4, 1]. Although fundamental, this approach inherently struggles with covariate shift and tasks that require memory beyond current observation [3, 11]. Many contemporary methods, despite advances, effectively operate within similar constraints or rely on limited observation histories. For instance, the Action Chunking Transformer (ACT) [5, 6], leveraging the Transformer architecture [18], predicts chunks of actions $a_{t:t+K-1}$ conditioned on present observation and potentially a latent variable from a CVAE. Although action chunking improves temporal smoothness and reduces the effective horizon [19, 20], its temporal modeling is largely confined to short-term dependencies implicitly captured through the time aggregation of chunks while reasoning, potentially failing when resolving ambiguities requires longer context [5]. Similarly, Diffusion Policy [7, 8, 9, 10], while adept at capturing complex, multimodal action distributions [7, 21], commonly conditions the diffusion process on the present observation or a short history, limiting its capacity for tasks requiring long-term memory [7, 21]. While extensions like Diff-Control [21] introduce forms of statefulness, they differ fundamentally from MTIL’s direct use of a recurrent SSM state to encode the full task history. Other techniques, including Implicit BC [22] and Energy-Based Models [23], also often operate primarily on the current state.
2.2 Temporal and Sequential Imitation Learning
The inadequacy of the strict Markov assumption has long motivated efforts to incorporate temporal context. Early explorations employed Recurrent Neural Networks (RNNs) like LSTMs [1, 4, 12, 24]. However, these architectures face challenges with long-term dependencies due to vanishing gradients [24]. Furthermore, practical implementations often resorted to fixed history windows and periodic state resets (e.g., sequence lengths of 10-50 in Robomimic [12, 25]), precluding the capture of full trajectory history. More recently, Transformer-based models (e.g., BeT [26], RT-1 [27, 28], OPTIMUS [29], ICRT [30], Baku [27, 28], MDT [31]) have become prominent, utilizing self-attention to model sequence correlations. Yet, the $O(L^{2})$ computational complexity of attention imposes practical limits on the size of context window  [1, 24], hindering their ability to efficiently process and encode entire long trajectories in the manner afforded by linear-complexity SSMs. Distinct strategies for managing long horizons involve temporal abstraction. Hierarchical Imitation Learning [14, 15, 32, 33, 34] and Skill Chaining [14, 35, 36] decompose tasks, learning policies over skills or sub-goals. Waypoint-based methods like AWE [19] or primitive-based approaches like PRIME [15] operate at higher levels of abstraction. While effective, these approaches fundamentally differ from MTIL, which aims to directly model the complete low-level observation-action sequence history, potentially offering robustness against issues like error propagation in skill chaining [14, 36].
2.3 State Space Models (SSMs) and Mamba in Robotics
State Space Models (SSMs) represent a compelling paradigm for sequence modeling, defined by their recurrent hidden state dynamics [16, 17, 37, 38, 39, 40, 41]. Mamba [16] marked a significant advancement, introducing input-dependent parameters ($\mathbf{A},\mathbf{B},\mathbf{C},\Delta$) via a selective scan mechanism. This allows Mamba models to dynamically focus on relevant sequence information while maintaining the linear time complexity characteristic of SSMs, synergizes the capacity for long-range dependency modeling, akin to Transformers, with the efficient recurrent updates reminiscent of RNNs, yet sidesteps the quadratic scaling bottlenecks of the former [1, 24] and the gradient propagation issues of the latter [24]. achieving strong empirical results [16].The robotics community has begun investigating Mamba’s potential [16, 17, 42, 43]. For instance, MaIL [17] employed Mamba as an imitation learning backbone, showing promise particularly in low-data regimes [16]. Mamba Policy [44, 45] integrated Mamba structures within diffusion models to enhance efficiency, while X-IL [43] explored Mamba within a modular IL framework. While these works adeptly leverage Mamba’s sequence processing power, MTIL distinguishes itself through its core premise: harnessing the step-updated recurrent state $h_{t}$ as an explicit, dynamically built representation of the entire observation history.This approach, tightly coupled with its sequential training methodology, directly overcomes the temporal ambiguity challenges inherent in Markovian assumptions common in imitation learning.
3 Mamba Temporal Imitation Learning (MTIL)
We introduce Mamba Temporal Imitation Learning (MTIL), a novel imitation learning framework designed to overcome the limitations of the Markov assumption by leveraging the full history of observations encoded within the recurrent state of an advanced State Space Model (SSM) architecture.
Figure 1: Overview of the Mamba Temporal Imitation Learning (MTIL) architecture. Multi-modal inputs (images via DINOv2, state) are fused and processed by sequential Mamba-2 blocks, updating the recurrent state $h_{t}$ which encodes history. At each step $t$ across the entire trajectory, MTIL predicts an action chunk $\hat{a}_{t:t+K-1}$ (current plus $K-1$ future steps). This is supervised via L2 loss against ground truth actions $a_{t:t+K-1}$ from the demonstration (using last action for padding when near trajectory end). The historical context embedded in $h_{t}$ enables temporally coherent, long-horizon action generation.
3.1 Background and Motivation
Standard imitation learning often assumes a Markov Decision Process (MDP), learning reactive policies $\pi(a_{t}|o_{t})$ via Behavioral Cloning. However, observational ambiguity fundamentally renders many sequential tasks as Partially Observable MDPs (POMDPs) [1], where the optimal policy necessitates conditioning on the full history $H_{t}=(o_{1},a_{1},...,o_{t})$. Theoretically, this history is captured by the belief state $b_{t}=P(s_{t}|H_{t})$, dictating the optimal policy $\pi^{*}(a_{t}|b_{t})$ [46].
Directly computing or representing the belief state $b_{t}$ is generally intractable. This motivates learning a compressed history representation $h_{t}\approx b_{t}$ using recurrent models. State Space Models (SSMs) like Mamba [16] offer a particularly compelling approach, providing a structured recurrent update $h_{t}=f(h_{t-1},x_{t})$ (where $x_{t}$ encodes $o_{t}$) with linear time complexity $O(L)$. This efficiency is critical for tractably encoding the long sequences required for full history representation, overcoming limitations of prior architectures [47].
Our motivation for MTIL stems from leveraging Mamba’s state $h_{t}$ as this potent, efficiently computed representation of the full history. By conditioning actions on both the current observation $o_{t}$ and this history-infused state $h_{t}$, MTIL learns a non-Markovian policy:
$\hat{a}_{t:t+K-1}\approx\pi(o_{t},h_{t})$
thereby directly addressing the core challenge of decision-making under ambiguity in POMDP-structured imitation learning by effectively utilizing the entire history.
3.2 Leveraging Full Trajectory History with Mamba-2
MTIL employs Mamba-2 [48], an advanced structured State Space Model (SSM) notable for its refined selective mechanism and efficiency [16]. Improving upon Mamba, Mamba-2 enhances hardware utilization and clarifies theoretical links to attention while retaining dynamic context adaptation via input-dependent parameters [48]. Its core lies in the discretized SSM recurrence governing the hidden state $h_{t}\in\mathbb{R}^{N}$ evolution based on input $x_{t}$ (derived from observation $o_{t}$):
$h_{t}=\bar{\mathbf{A}}_{t}h_{t-1}+\bar{\mathbf{B}}_{t}x_{t}$
$y_{t}=\mathbf{C}_{t}h_{t}$
Crucially, the input-dependent parameters $(\Delta_{t},\bar{\mathbf{B}}_{t},\mathbf{C}_{t})=f(x_{t})$ enable selective state dynamics, allowing the model to dynamically modulate information flow and preserve salient historical context. This selective mechanism, combined with inherent linear-time complexity $O(L)$, facilitates learning from complete trajectory histories—a significant advantage over quadratic-complexity $O(L^{2})$ attention mechanisms or memory-limited RNNs for long sequences. The resulting state $h_{t}$ acts as a dynamic summary of the salient history $(x_{1},...,x_{t-1})$, furnishing the requisite context even when the instantaneous observation $x_{t}$ is ambiguous. The MTIL policy leverages this directly:
$\pi(\hat{a}_{t:t+K-1}|x_{t},h_{t})$
By conditioning predictions $\hat{a}_{t:t+K-1}$ on both the current input $x_{t}$ and the comprehensive historical summary encoded in $h_{t}$, MTIL effectively transcends the limitations inherent in Markovian approaches, enabling sequential decision-making grounded in the full trajectory context.
Algorithm 1 MTIL Training (Sequential Step-based)
0:  Expert trajectories $\mathcal{D}=\{\tau_{i}\}$, $\tau_{i}=(o_{1},a_{1},...,o_{T_{i}},a_{T_{i}})$, MTIL Policy $\pi_{\theta}$, initialized parameters $\theta$, Loss $\mathcal{L}$ (MSE), Chunk size $k$, Optimizer $Opt$
1:  Initialize policy parameters $\theta$
2:  for each training epoch do
3:     for each trajectory $\tau_{i}\in\mathcal{D}$ do
4:        Initialize hidden state $h_{0}$, trajectory loss $\mathcal{L}_{traj}\leftarrow 0$
5:        for $t=0$ to $T_{i}-1$ do
6:           $x_{t}=\text{Encoder}(o_{t})$
7:           Predict $(\hat{a}_{t:t+k-1},h_{t})=\pi_{\theta}.\text{step}(x_{t},h_{t-1})$
8:           Get $a_{t:t+k-1}$ from $\tau_{i}$
9:           Calculate step loss: $\mathcal{L}_{t}=\mathcal{L}(\hat{a}_{t:t+k-1},a_{t:t+k-1})$
10:           Accumulate loss: $\mathcal{L}_{traj}+=\mathcal{L}_{t}$
11:        end for
12:        $Opt.\text{step}()$Update $\theta$}
13:     end for
14:  end for
15:  return  Trained policy $\pi_{\theta}$.
Algorithm 2 MTIL Inference (with Action Chunking and Temporal Aggregation)
0:  Trained policy $\pi_{\theta}$, Initial observation $o_{0}$, Chunk size $k$, Max steps $T_{max}$, Exponential aggregation weights $W$
1:  Initialize hidden state $h_{0}$, prediction buffer $B$.
2:  for $t=0$ to $T_{max}-1$ do
3:     $x_{t}=\text{Encoder}(o_{t})$
4:     Predict $(\hat{a}_{t:t+k-1},h_{t})=\pi_{\theta}.\text{step}(x_{t},h_{t-1})$
5:     Store prediction $\hat{a}_{t:t+k-1}$ in buffer $B$
6:     Aggregate Action for step $t$:
7:      Get predictions for step $t$ from $B$: $P_{t}=\{\hat{a}_{j:j+k-1}[t-j]\mid j\leq t<j+k\text{ and }\hat{a}_{j:j+k-1}\in B\}$
8:      Compute final action: $a_{t}^{\text{final}}=\text{WeightedAverage}(P_{t},W)$
9:     Execute action $a_{t}^{\text{final}}$
10:  end for
3.3 MTIL Training and Inference
MTIL enables imitation learning across complete expert trajectories, utilizing the architecture outlined in Figure 1. Distinctively, MTIL employs a sequential training procedure (Algorithm 1). This step-wise paradigm, leveraging Mamba’s recurrent ‘step‘ function, is fundamental to efficiently encoding arbitrarily long trajectory histories from high-dimensional observations (e.g., images) within feasible memory constraints—a key departure from parallel window-based approaches.At each timestep $t$ along the trajectory (length $T$), the policy receives the observation embedding $x_{t}$, updates its history-encoding hidden state from $h_{t-1}$ to $h_{t}$, and predicts a chunk of $K$ actions $\hat{a}_{t:t+K-1}$, covering the current and $K-1$ future steps. Learning proceeds by minimizing the Mean Squared Error (MSE) between this predicted chunk $\hat{a}_{t:t+K-1}$ and the ground truth actions $a_{t:t+K-1}$ sourced from the demonstration data. Notably, when $t+K-1$ exceeds the trajectory length $T$, the ground truth chunk is padded by repeating the final demonstrated action $a_{T}$. During inference (Algorithm 2), the trained policy operates autoregressively. It uses the same ‘step‘ function to update its hidden state $h_{t}$ based on the current observation and predict the action chunk $\hat{a}_{t:t+K-1}$. For enhanced stability and smoothness in execution, temporal aggregation strategies [19, 20, 5] are commonly applied. This involves computing the final action for time $t$ by applying a weighted average over multiple predictions made for this specific timestep but generated at different preceding steps.
4 Experimental Results
We conducted extensive experiments to evaluate the performance of MTIL across various benchmarks and real-world scenarios.See more details in Appendix A
4.1 ACT benchmark
Table 1: Success rates(%) on ACT.
Cube Bimanual
Method Transfer Insertion
ACT [5] 90 50
MTIL (10-step) 92 56
MTIL (Full) 100 84
MTIL’s efficiency and performance were evaluated on ACT [5] simulation tasks. Table 1 presents the final success rates, while Figure 2 shows learning curves and a backbone ablation. Compared to ACT [5], MTIL leveraging full trajectory history (approx.400 steps in these tasks) achieves substantially higher success rates (Table 1: 100%/84% vs.90%/50% for ACT on Cube Transfer/Bimanual Insertion) and markedly faster convergence (Figure 2a). We also evaluated MTIL variants using shorter fixed history windows (10, 20, and 50 steps),and their final performance did not show statistically significant differences from each other (10-step results: 92%/56%), confirming that such short-history modeling is clearly inferior to utilizing the full-history available. Furthermore, to isolate the benefits of the full-history approach from visual feature quality, we conducted an ablation comparing MTIL using DINOv2 [49] against ResNet18 [50] (which is used by ACT). While DINOv2 offers stronger features—a factor contributing to its selection, particularly for real-world robustness—the results (Figure 2b) combined with the short-versus-full history comparison underscore that the primary performance leap stems from MTIL’s core mechanism: encoding comprehensive trajectory context.
(a) Learning Curves
(b) Backbone Comparison
Figure 2: ACT benchmark performance.(a) Learning curves (Cube Transfer left, Bimanual Insertion right). (b) Backbone comparison (MTIL Full-History, DINOv2 vs. ResNet18).
4.2 LIBERO Benchmark
[…]
Table 2: Lifelong Learning Performance on LIBERO (EWC Strategy).
LIBERO-LONG LIBERO-SPATIAL
Policy Arch. FWT($\uparrow$) NBT($\downarrow$) AUC($\uparrow$) FWT($\uparrow$) NBT($\downarrow$) AUC($\uparrow$)
ResNet-RNN 0.02 $\pm$ 0.00 0.04 $\pm$ 0.01 0.00 $\pm$ 0.00 0.14 $\pm$ 0.02 0.23 $\pm$ 0.02 0.03 $\pm$ 0.00
ResNet-T 0.13 $\pm$ 0.02 0.22 $\pm$ 0.03 0.03 $\pm$ 0.00 0.23 $\pm$ 0.01 0.33 $\pm$ 0.01 0.06 $\pm$ 0.01
ResNet-M (10-step) 0.14 $\pm$ 0.02 0.20 $\pm$ 0.03 0.03 $\pm$ 0.00 0.24 $\pm$ 0.01 0.30 $\pm$ 0.02 0.06 $\pm$ 0.01
ResNet-M (Full) 0.22 $\pm$ 0.03 0.08 $\pm$ 0.02 0.08 $\pm$ 0.02 0.28 $\pm$ 0.02 0.17 $\pm$ 0.02 0.05 $\pm$ 0.01
ViT-T 0.05 $\pm$ 0.02 0.09 $\pm$ 0.03 0.01 $\pm$ 0.00 0.32 $\pm$ 0.01 0.48 $\pm$ 0.03 0.06 $\pm$ 0.01
ViT-M (10-step) 0.06 $\pm$ 0.02 0.10 $\pm$ 0.03 0.01 $\pm$ 0.00 0.33 $\pm$ 0.01 0.45 $\pm$ 0.03 0.06 $\pm$ 0.01
ViT-M (Full) 0.19 $\pm$ 0.04 0.05 $\pm$ 0.01 0.10 $\pm$ 0.03 0.35 $\pm$ 0.02 0.15 $\pm$ 0.03 0.10 $\pm$ 0.01
LIBERO-OBJECT LIBERO-GOAL
Policy Arch. FWT($\uparrow$) NBT($\downarrow$) AUC($\uparrow$) FWT($\uparrow$) NBT($\downarrow$) AUC($\uparrow$)
ResNet-RNN 0.17 $\pm$ 0.04 0.23 $\pm$ 0.04 0.06 $\pm$ 0.01 0.16 $\pm$ 0.01 0.22 $\pm$ 0.01 0.06 $\pm$ 0.01
ResNet-T 0.56 $\pm$ 0.03 0.69 $\pm$ 0.02 0.16 $\pm$ 0.02 0.32 $\pm$ 0.04 0.45 $\pm$ 0.04 0.07 $\pm$ 0.01
ResNet-M (10-step) 0.50 $\pm$ 0.03 0.39 $\pm$ 0.03 0.15 $\pm$ 0.02 0.31 $\pm$ 0.04 0.42 $\pm$ 0.04 0.07 $\pm$ 0.01
ResNet-M (Full) 0.55 $\pm$ 0.03 0.36 $\pm$ 0.03 0.17 $\pm$ 0.01 0.30 $\pm$ 0.03 0.11 $\pm$ 0.04 0.10 $\pm$ 0.01
ViT-T 0.57 $\pm$ 0.03 0.64 $\pm$ 0.03 0.23 $\pm$ 0.00 0.32 $\pm$ 0.04 0.48 $\pm$ 0.03 0.07 $\pm$ 0.01
ViT-M (10-step) 0.56 $\pm$ 0.03 0.60 $\pm$ 0.03 0.22 $\pm$ 0.01 0.33 $\pm$ 0.04 0.45 $\pm$ 0.03 0.08 $\pm$ 0.01
ViT-M (Full) 0.58 $\pm$ 0.03 0.18 $\pm$ 0.04 0.25 $\pm$ 0.01 0.34 $\pm$ 0.04 0.10 $\pm$ 0.03 0.11 $\pm$ 0.01
FWT($\uparrow$): Forward Transfer, NBT($\downarrow$): Backward Transfer, AUC($\uparrow$): Area Under Curve. EWC strategy results averaged over 3 seeds (100, 200, 300) at 50 epochs. Baselines from [51]. Short-history (10-step,similar performance for 20/50 steps) and full-history result shown.
4.3 Robomimic (Low-Dimensional State)
To assess MTIL’s ability to model complex dynamics from proprioceptive states alone, we evaluated it on low-dimensional Robomimic tasks [12], comparing against prior methods including Diffusion Policy [7], Even without visual input, MTIL leveraging full history consistently surpasses previous state-of-the-art methods across the evaluated tasks.(Table 3).
Table 3: Success Rates on Robomimic Low-Dimensional Tasks.
Lift Can Square Transport Tool Hang
Method ph mh ph mh ph mh ph mh ph mh
LSTM-GMM [12] 1.00 0.93 1.00 0.81 0.95 0.59 0.76 0.20 0.67 0.31
IBC [22] 0.79 0.15 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.00
BET [26] 1.00 0.99 1.00 0.90 0.76 0.43 0.38 0.06 0.58 0.20
DiffusionPolicy-C [7] 1.00 0.97 1.00 0.96 0.97 0.82 0.94 0.46 0.50 0.30
DiffusionPolicy-T [7] 1.00 1.00 1.00 0.94 1.00 0.81 1.00 0.35 1.00 0.87
MTIL (10-step) 1.00 0.94 1.00 0.83 0.96 0.61 0.78 0.22 0.70 0.33
MTIL (Full History) 1.00 1.00 1.00 0.96 0.98 0.83 1.00 0.48 1.00 0.89
An MTIL variant using only a 10-step history performs comparably, albeit slightly better, than LSTM-GMM [12], indicating limited gains from short recurrent contexts here. The full-history configuration, however, demonstrates the strength of Mamba’s architecture within MTIL for capturing intricate state transition dynamics and temporal dependencies inherent in manipulation, proving effective beyond visuomotor control.
4.4 Real-World Dual-Arm Tasks
Figure 3: Dual UR3 experimental setup with four cameras (Top: Kinect; Side: D435i; Wrists: D405) and custom grippers.
To validate MTIL in complex physical environments, we designed challenging tasks on a dual UR3 platform equipped with custom 2-finger grippers and four cameras providing multi-view observations (Figure 3). We compare MTIL (using DINOv2 backbone and full history) against ACT trained on identical demonstration data (100 demos per task).
Sequential Insertion Task.
We designed this task (visualized in Figure 4) specifically to challenge Markovian policies by requiring long-term memory, a scenario where methods like ACT often fail. The four stages involve: (1) Left arm grasps Tube1, (2) Left passes Tube1 to Right arm, (3) Right arm inserts Tube1 into Tube2, (4) Right arm inserts Tube1 into Tube3. Critically, executing Stage 3 correctly necessitates recalling the completion of previous stages, as intermediate observations can be ambiguous. Table 4 details the stage-wise success rates. MTIL, leveraging its full history state, successfully completes the entire sequence with high probability. In stark contrast, ACT, reliant on immediate context, is confounded by the temporal ambiguity; it frequently attempts Stage 4 directly after Stage 2, failing to execute the required sequence correctly and resulting in zero success for completing Stage 3, Stage 4, and the overall task. This outcome underscores the limitations of short-history approaches and validates the imperative of encoding complete history for reliably executing temporally complex manipulation sequences.
Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
ACT 80 64 0 0 0
MTIL 94 80 62 54 54
Figure 4: Figure & Table 4: Sequential Insertion task stages (left) and success rates (%, right). MTIL successfully completes the sequence while ACT fails due to temporal ambiguity between stages 3 and 4.
Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
ACT 80 64 50 32 32
MTIL 94 80 74 62 62
Figure 5: Figure & Table 5: Coordinated Pouring task stages (left) and success rates (%, right). MTIL achieves higher success and smoother execution compared to ACT.
Coordinated Pouring Task.
This task (Figure 5) assesses precise bimanual coordination over a longer sequence: (1) Left arm grasps Tube1, (2) Left passes Tube1 to Right arm, (3) Left arm grasps Tube2, (4) Right arm pours water from Tube1 into Tube2. While less susceptible to the specific ambiguity of the insertion task, it still requires accurate, temporally coordinated actions. Table 5 shows that although both methods achieve non-zero success, MTIL consistently outperforms ACT across the stages, resulting in a higher overall success rate and exhibiting notably smoother execution trajectories.
5 Conclusion
The trajectory of intelligence is intrinsically linked to the capacity for memory – the ability to weave the tapestry of past experiences into the fabric of present action. This work confronts a central limitation in contemporary imitation learning: the prevalent reliance on the Markovian assumption, which often reduces complex sequential behaviors to mere reactions to the immediate sensory world, neglecting the crucial dimension of history. We introduced Mamba Temporal Imitation Learning (MTIL), a paradigm shift that embraces the power of memory by leveraging the recurrent state dynamics inherent within the Mamba architecture. MTIL demonstrates that by encoding the full history of observations into a compressed, evolving state representation, robots can transcend the confines of the present moment. This comprehensive temporal context allows MTIL to effectively disambiguate perception and unlock the execution of intricate, state-dependent sequential tasks previously challenging for established methods like ACT and Diffusion Policy, as validated through extensive experiments across diverse simulated and real-world benchmarks. Our findings not only showcase the significant performance and efficiency gains afforded by MTIL but, more profoundly, underscore the essential role of history in bridging the gap between perception and intelligent, temporally coherent action in robotics. By demonstrating the efficacy of State Space Models in capturing the long flow of time, this work illuminates a promising pathway towards building robotic agents capable of deeper understanding and more sophisticated interaction with the world.
Limitations
While MTIL demonstrates significant advantages, we acknowledge certain limitations. Firstly, our current training strategy relies on sequential processing via Mamba’s ‘step‘ function(as detailed in Section 3.3). This is necessitated by the significant GPU memory demands of processing entire long sequences of high-dimensional observations concurrently, a challenge for current hardware. Unlike methods operating under the Markov assumption which might permit data shuffling, MTIL must preserve temporal order, inherently limiting parallelization compared to approaches like windowed Transformers and resulting in slower wall-clock training times per epoch—a trade-off for enabling full history modeling. Exploring techniques like gradient checkpointing or specialized hardware optimizations presents an avenue to mitigate this training speed limitation. Secondly, a potential failure mode, common to many imitation learning approaches, stems from distributional shift between training demonstrations and deployment scenarios. MTIL’s generalization capability to substantially novel environments or configurations may be limited, potentially leading to performance degradation when facing significant train-test mismatch.Lastly, while Mamba’s state aims to compress history effectively, its fidelity over ultra-long sequences (e.g., tens of thousands of steps) remains to be rigorously verified. Furthermore, performance can exhibit sensitivity to hyperparameter choices within the Mamba architecture and training regime, necessitating careful tuning. Addressing these aspects, alongside further theoretical investigation into the state representation limits, remains important future work.
Acknowledgments
This work is supported by the Joint Funds of the National Natural Science Foundation of China (Grant No. U [^2] A20208), the Natural Science Foundation Innovation Group Project of Hubei Province (Grant No. 2022CFA018), and the Key Research and Development Program of Guangdong Province (Grant No. 2022B0202010001-2).
Appendix A Appendix
A.1 ACT Benchmark Experimental Setup
Tasks and Data.
We evaluated MTIL on the two simulated tasks introduced in the ACT paper [5]: ”Cube Transfer” and ”Bimanual Insertion”. For both tasks, we used a dataset of 100 expert demonstration trajectories, automatically generated using scripted policies provided in the original ACT repository. The typical trajectory length for these tasks is around 400 steps.
MTIL Configuration.
The main MTIL model (”MTIL (Full)”) used full history encoding, processing the entire trajectory sequentially during training as described in Algorithm 1. We used an action chunk prediction size of $K=50$. The model was trained using the AdamW optimizer [53] with an initial learning rate of $2\times 10^{-4}$. A cosine learning rate decay schedule was applied over the training epochs. For the short-history ablation (”MTIL (10-step Hist.)”), the same architecture and training setup were used, but the hidden state was reset every 10 steps, limiting the effective history length.
Backbones and Feature Processing.
We experimented with two visual backbones to compare performance and isolate the effect of history modeling:
  • ResNet18 (for fair comparison with ACT): Input RGB images (480x640x3) from multiple cameras were processed by separate ResNet18 [50] instances. Following the ACT pipeline, we extracted feature maps from an intermediate layer, resulting in 15x20x512 feature maps per camera. We applied learned positional encodings to these features. To enhance feature representation, we incorporated a Convolutional Block Attention Module (CBAM) [54] after the ResNet features, applying both channel and spatial attention mechanisms. The resulting features were flattened and projected using a linear layer to a common embedding dimension.
  • DINOv2 (ViT-L/14, for best performance): Input RGB images (480x640x3) were fed into a pre-trained DINOv2 model with a ViT-L/14 architecture [49]. The weights of the DINOv2 model were kept frozen during training. We extracted features from the output of the penultimate transformer block (specifically, ‘blocks[-4]‘ in common implementations), yielding a feature map of size approximately 34x45x1024 per camera (exact dimensions depend slightly on implementation details of patch embedding and image size). These high-dimensional features were then flattened and projected via a linear layer to the common embedding dimension.
Multi-Modal Fusion.
The projected visual features from all cameras were concatenated. Low-dimensional robot state information (e.g., joint positions, gripper width/status) was also concatenated into a vector. To effectively integrate these modalities, we employed a cross-modal attention mechanism: the concatenated visual features served as query, while the low-dimensional state vector served as the keys and values. The output of this attention mechanism was then passed through a final linear layer to project it to the input dimension ($D$) required by the Mamba-2 blocks.
Mamba-2 Architecture.
The core sequence modeling was performed by a stack of Mamba-2 blocks. Key hyperparameters such as the model dimension ($d_{model}$), state expansion factor, state dimension ($d_{state}$), and the number of layers were tuned based on validation performance. Specific values are provided in the codebase and Table 6.
A.2 LIBERO Benchmark Experimental Setup
Goal and Setup.
The experiments on LIBERO [51] aimed to evaluate MTIL’s performance in a challenging lifelong learning setting, focusing on knowledge transfer and catastrophic forgetting. To ensure fair comparison with the baselines reported in [51], we adhered closely to their experimental protocol.
Backbones and Prediction.
We used the standard ResNet and ViT [55] backbones provided within the official LIBERO benchmark codebase. We did *not* use the DINOv2 backbone in these experiments to maintain comparability. Furthermore, following the LIBERO setup for RNN/Transformer baselines, MTIL was configured to predict only the action for the next single step ($K=1$), rather than action chunks.
Language Encoder and Fusion.
Task descriptions were encoded using the standard BERT [56] encoder, consistent with the LIBERO baselines. For fusion, simple concatenation of visual features, state information, and language embeddings followed by linear projection was used, matching the baseline setup. No cross-modal attention mechanism was employed here.
Lifelong Learning Strategy: EWC.
We employed the Elastic Weight Consolidation (EWC) [52] strategy to mitigate catastrophic forgetting as tasks were learned sequentially. EWC works by calculating the Fisher Information Matrix (FIM) for the parameters of the model with respect to previously learned tasks. During training on a new task, a quadratic penalty term is added to the loss function, discouraging changes to parameters deemed important for past tasks (i.e., those with high Fisher information). Rationale for EWC with MTIL: We selected EWC not only because it is a standard lifelong learning technique used in the benchmark, but also due to its theoretical synergy with MTIL’s state-based history encoding. The Mamba-2 blocks contain parameters that govern the selective state updates ($\Delta_{t},\mathbf{B}_{t},\mathbf{C}_{t}$) and state transitions ($\mathbf{A}_{t}$). These parameters are crucial for capturing the temporal dynamics and historical context learned from demonstrations. By applying EWC, we hypothesize that the FIM helps identify and protect those parameters within the Mamba architecture that are most critical for encoding the temporal patterns learned in previous tasks. This allows MTIL to retain its ability to leverage history effectively for older tasks (leading to low NBT) while still adapting its state dynamics to accommodate new tasks (enabling high FWT).
Metrics Definition.
We report the standard LIBERO lifelong learning metrics, averaged over 3 random seeds after training on all tasks in a suite for 50 epochs using the EWC strategy:
  • FWT (Forward Transfer, $\uparrow$): Measures the influence of having learned tasks $11,...,i-$ 1 on the initial performance of learning task $i$, compared to learning task $i$ from scratch. It is calculated as:
    $FWT=\frac{1}{N-1}\sum_{i=2}^{N}(A_{i-1,i}-A_{base,i})$
    where $A_{i-1,i}$ is the accuracy on task $i$ evaluated *immediately after* training on task $i-1$ (before training on task $i$), and $A_{base,i}$ is the accuracy on task $i$ when trained in isolation (from scratch). Higher values indicate better positive transfer. (Note: LIBERO paper definition might vary slightly, we follow the standard interpretation of immediate transfer).
  • NBT (Negative Backward Transfer, $\downarrow$): Measures the average drop in performance on previously learned tasks after learning subsequent tasks (i.e., forgetting). It is calculated as:
    $NBT=\frac{1}{N-1}\sum_{i=1}^{N-1}(\max(0,A_{i,i}-A_{N,i}))$
    where $A_{i,i}$ is the final accuracy on task $i$ after training on task $i$, and $A_{N,i}$ is the accuracy on task $i$ after training on all $N$ tasks. Lower values indicate less forgetting.
  • AUC (Area Under the learning Curve, $\uparrow$): Represents the overall performance across all tasks throughout the lifelong learning process. It is typically calculated as the average accuracy across all tasks, evaluated after the model has finished training on the final task $N$:
    $AUC=\frac{1}{N}\sum_{i=1}^{N}A_{N,i}$
    Higher values indicate better overall performance retention and learning.
A.3 Robomimic Benchmark Experimental Setup
Goal and Setup.
The experiments on Robomimic [12] aimed to evaluate MTIL’s ability to model complex manipulation skills using only low-dimensional proprioceptive state information (e.g., joint positions, velocities, end-effector pose, gripper status), without any visual input.
Input and Settings.
We used the standard low-dimensional state datasets provided by Robomimic for tasks like Lift, Can, Square, Transport, and Tool Hang, considering both proficient-human (ph) and multi-human (mh) data sources. To ensure fair comparison, we followed the evaluation protocol and data splits used in prior work, particularly matching the setup for the LSTM-GMM baseline [12] where applicable.
Key Difference and Action Output.
The primary difference between MTIL and the LSTM-GMM baseline lies in the history modeling. While LSTM-GMM uses LSTMs, often trained on shorter sequences (e.g., 10-50 steps) in practice [12], MTIL (Full History) processes the entire demonstration sequence using its Mamba-based recurrent state. To isolate the effect of the sequence modeling backbone, we configured MTIL to use a Gaussian Mixture Model (GMM) head for predicting the action distribution at each step, mirroring the output structure of LSTM-GMM. The number of mixture components was treated as a hyperparameter.
MTIL Architecture.
A stack of Mamba-2 blocks was used to process the sequence of low-dimensional state vectors. The specific hyperparameters (e.g., $d_{model}$, $d_{state}$, number of layers) were optimized for these tasks (See details in Table 6).
A.4 Real-World Experimental Setup
Figure 6: Details of the experimental hardware, showing the dual-arm robot setup (UR3) with the task workspace, and the specific custom gripper, wrist camera(D405), Kinect camera and RealSense D435 utilized.
Platform.
Experiments were conducted on a physical setup consisting of two Universal Robots UR3 arms mounted on a table. Each arm was equipped with a custom 3D-printed 2-finger parallel gripper actuated by a Dynamixel servo motor. The sensory input comprised multi-view images from four cameras: one Intel RealSense D435i providing a side view, one Microsoft Azure Kinect providing a top-down view, and two Intel RealSense D405 cameras mounted on the wrists of the robots for close-up views (see Figure  6).
MTIL Configuration.
For the real-world tasks (Sequential Insertion, Coordinated Pouring), we used the full MTIL configuration designed for robust performance: full history encoding via sequential training (Algorithm 1), the DINOv2 (ViT-L/14) backbone with frozen weights for visual feature extraction, and prediction of action chunks with $K=50$. Multi-modal fusion using cross-attention between visual features and low-dimensional state (joint positions, gripper state) was employed, identical to the setup described in Appendix A.1. Specific Mamba-2 architecture parameters were consistent with those used in the ACT benchmark simulation experiments with DINOv2.
Data Collection.
For each of the two tasks, we collected 100 expert demonstrations using a custom teleoperation system that use spacemouse. The system allowed a human operator to control the dual arms simultaneously via input devices, recording the robot states and camera images at each timestep.The trajectory length for the Sequential Insertion task is around 500 steps while the trajectory length for the Coordinated Pouring task is around 1000 steps.
Inference Details: Chunking and Temporal Aggregation.
During policy execution on the real robot (inference time, Algorithm 2), we applied both action chunking (predicting $K=50$ steps ahead) and temporal aggregation. Temporal aggregation was implemented using a weighted averaging scheme over overlapping predicted chunks. Specifically, the action $a_{t}^{\text{final}}$ executed at time $t$ was computed as a weighted average of predictions made for time $t$ from chunks initiated at previous times $j$ (where $j\leq t<j+K$). We used exponentially decaying weights $W=\{w_{0},...,w_{K-1}\}$ where $w_{k}\propto\gamma^{k}$ for some decay factor $\gamma<1$, giving higher importance to predictions made closer to the current time step (i.e., smaller $k=t-j$). Rationale: Applying chunking and temporal aggregation is a common practice in methods like ACT [5] and has been shown to significantly improve the smoothness and stability of executed robot motions by reducing jitter [19, 20]. We employed this technique primarily for two reasons: (1) To ensure a fair comparison with ACT, which also utilizes this strategy. (2) To achieve better practical performance and robustness during real-world execution, as smoother actions are generally preferable. While MTIL’s inherent temporal modeling capability provides a strong foundation for coherent action sequences, the temporal aggregation acts as an additional filter that further enhances stability, leading to more reliable task execution.
A.5 Hyperparameter Summary
Table 6: Key MTIL Hyperparameters Across Benchmarks.
Parameter ACT Sim LIBERO Robomimic Real World
Backbone ResNet18 / DINOv2 ResNet / ViT N/A (State) DINOv2
History Full / 10-step Full / 10-step Full / 10-step Full
Chunk Size (K) 50 1 1 50
Mamba $d_{model}$ 2048 256 64 2048
Mamba $d_{state}$ 512 64 16 512
Mamba Layers 4 1 4 4
Optimizer AdamW AdamW AdamW AdamW
Initial LR 2e-4 1e-4 1e-4 2e-4
LR Schedule Cosine Cosine Cosine Cosine
Cross-Attention Yes No N/A Yes
Output Head Linear Linear GMM Linear
Temporal Aggregation Yes No No Yes

### v3

@@
 Abstract
-Standard imitation learning (IL) methods have achieved considerable success in robotics, yet often rely on the Markov assumption, limiting their applicability to tasks where historical context is crucial for disambiguating current observations.
-This limitation hinders performance in long-horizon sequential manipulation tasks where the correct action depends on past events not fully captured by the current state.
-To address this fundamental challenge, we introduce Mamba Temporal Imitation Learning (MTIL), a novel approach that leverages the recurrent state dynamics inherent in State Space Models (SSMs), specifically the Mamba architecture.
-MTIL encodes the entire trajectory history into a compressed hidden state, conditioning action predictions on this comprehensive temporal context alongside current multi-modal observations.
-Through extensive experiments on simulated benchmarks (ALOHA, Robomimic, LIBERO) and on challenging real-world sequential manipulation tasks that requiring long-term temporal reasoning, MTIL significantly outperforms state-of-the-art methods like ACT and Diffusion Policy.
-Our findings affirm the necessity of full temporal context for robust sequential decision-making and validate MTIL as a powerful approach that transcends the inherent limitations of Markovian imitation learning.
-Our implementation is available at MTIL
-Keywords: Imitation Learning, Robot Manipulation, Sequence Modeling, State Space Models
-1 Introduction
+Standard imitation learning (IL) methods have achieved considerable success in robotics, yet often rely on the Markov assumption, which falters in long-horizon tasks where history is crucial for resolving perceptual ambiguity.
+This limitation stems not only from a conceptual gap but also from a fundamental computational barrier: prevailing architectures like Transformers are often constrained by quadratic complexity, rendering the processing of long, high-dimensional observation sequences infeasible.
+To overcome this dual challenge, we introduce Mamba Temporal Imitation Learning (MTIL).
+Our approach represents a new paradigm for robotic learning, which we frame as a practical synthesis of World Model and Dynamical System concepts.
+By leveraging the linear-time recurrent dynamics of State Space Models (SSMs), MTIL learns an implicit, action-oriented world model that efficiently encodes the entire trajectory history into a compressed, evolving state.
+This allows the policy to be conditioned on a comprehensive temporal context, transcending the confines of Markovian approaches.
+Through extensive experiments on simulated benchmarks (ACT, Robomimic, LIBERO) and on challenging real-world tasks, MTIL demonstrates superior performance against SOTA methods like ACT and Diffusion Policy, particularly in resolving long-term temporal ambiguities.
+Our findings not only affirm the necessity of full temporal context but also validate MTIL as a powerful and a computationally feasible approach for learning long-horizon, non-Markovian behaviors from high-dimensional observations.
+Project code are available at https://github.com/yulinzhouZYL/MTIL
+I Introduction
 Imitation Learning (IL) has emerged as a powerful paradigm for teaching robots complex skills directly from expert demonstrations, bypassing the need for intricate reward engineering often required in reinforcement learning [1, 2, 3].
@@
 This temporal ambiguity signifies an underlying Partially Observable Markov Decision Process (POMDP), posing a critical challenge for standard IL methods in state-dependent tasks.
-Since ambiguous tasks manifest as POMDPs and human demonstrations are inherently non-Markovian, effective imitation necessitates history-aware policies.While the critical role of historical context has been increasingly recognized  [12], effectively and efficiently encoding comprehensive trajectory histories into actionable representations remains a significant challenge.
-Addressing this, we introduce Mamba Temporal Imitation Learning (MTIL).This approach is specifically designed to incorporate the complete observational history into decision-making by leveraging the unique properties of State Space Models (SSMs), particularly the recently developed Mamba architecture [16, 17].
+Since ambiguous tasks manifest as POMDPs and human demonstrations are inherently non-Markovian, effective imitation necessitates history-aware policies.
+While the critical role of historical context has been increasingly recognized [12], a fundamental barrier has remained: the computational infeasibility of processing long, high-dimensional observation histories with prevailing architectures like Transformers.
+Addressing this, we introduce Mamba Temporal Imitation Learning (MTIL).
+Our approach is specifically designed to incorporate the complete observational history into decision-making by leveraging the unique properties of State Space Models (SSMs), particularly the recently developed Mamba architecture [16, 17].
 Mamba’s recurrent structure allows it to maintain a compressed hidden state $h_{t}$ that theoretically encapsulates information from the entire preceding observation sequence $H_{t}=(o_{1},...,o_{t})$.
@@
   1. 1.
-    A novel Mamba-based IL architecture, MTIL, is proposed, specifically designed for full trajectory history encoding via its recurrent state to overcome limitations inherent in Markovian sequential task learning.
+    We propose MTIL, a novel architecture that is the first to leverage the linear-time recurrence of State Space Models to make full-trajectory imitation learning from high-dimensional visual data computationally feasible on commodity hardware, breaking the quadratic bottleneck of attention-based models.
   2. 2.
-    The fundamental challenge of temporal ambiguity in IL is addressed by incorporating comprehensive historical context, offering a promising pathway towards learning and executing long-horizon, state-dependent manipulation sequences previously intractable for methods reliant on limited context.
+    We provide a new theoretical framing for this approach, positing MTIL as learning an implicit dynamical system.
+This system’s evolving state acts as a continuous belief-state representation, offering a robust solution to the core problem of temporal ambiguity in partially observable environments.
   3. 3.
-    Extensive empirical validation is provided across diverse benchmarks (simulation and real-world), demonstrating significant performance improvements over existing state-of-the-art methods (including ACT, Diffusion Policy), particularly on tasks demanding temporal reasoning and long-term memory.
-2 Related Works
-2.1 Markovian and Short-History Imitation Learning
+    We provide extensive empirical validation demonstrating that MTIL significantly outperforms state-of-the-art methods, including ACT and Diffusion Policy.
+Furthermore, on tasks requiring long-term memory, MTIL also surpasses other full-history-capable baselines like Transformer-XL, validating the unique advantages of its underlying architecture.
+II Related Works
+II-A Markovian and Short-History Imitation Learning
 A cornerstone of imitation learning, Behavioral Cloning (BC), typically learns a Markovian policy $\pi(a_{t}|o_{t})$ via supervised learning [2, 3, 4, 1].
@@
 Other techniques, including Implicit BC [22] and Energy-Based Models [23], also often operate primarily on the current state.
-2.2 Temporal and Sequential Imitation Learning
+II-B Temporal and Sequential Imitation Learning
 The inadequacy of the strict Markov assumption has long motivated efforts to incorporate temporal context.
@@
 More recently, Transformer-based models (e.g., BeT [26], RT-1 [27, 28], OPTIMUS [29], ICRT [30], Baku [27, 28], MDT [31]) have become prominent, utilizing self-attention to model sequence correlations.
-Yet, the $O(L^{2})$ computational complexity of attention imposes practical limits on the size of context window  [1, 24], hindering their ability to efficiently process and encode entire long trajectories in the manner afforded by linear-complexity SSMs.
+However, the $O(L^{2})$ computational complexity of attention imposes practical limits on the size of the context window [1, 24], hindering their ability to efficiently process entire long trajectories.
+Even recurrent variants like Transformer-XL [32], while theoretically capable of processing long sequences, still rely on the computationally intensive attention mechanism.
 Distinct strategies for managing long horizons involve temporal abstraction.
@@
 While effective, these approaches fundamentally differ from MTIL, which aims to directly model the complete low-level observation-action sequence history, potentially offering robustness against issues like error propagation in skill chaining [14, 37].
-2.3 State Space Models (SSMs) and Mamba in Robotics
+II-C State Space Models (SSMs) and Mamba in Robotics
 State Space Models (SSMs) represent a compelling paradigm for sequence modeling, defined by their recurrent hidden state dynamics [16, 17, 38, 39, 40, 41, 42].
@@
 While these works adeptly leverage Mamba’s sequence processing power, MTIL distinguishes itself through its core premise: harnessing the step-updated recurrent state $h_{t}$ as an explicit, dynamically built representation of the entire observation history.This approach, tightly coupled with its sequential training methodology, directly overcomes the temporal ambiguity challenges inherent in Markovian assumptions common in imitation learning.
-3 Mamba Temporal Imitation Learning (MTIL)
+III Mamba Temporal Imitation Learning (MTIL)
 We introduce Mamba Temporal Imitation Learning (MTIL), a novel imitation learning framework designed to overcome the limitations of the Markov assumption by leveraging the full history of observations encoded within the recurrent state of an advanced State Space Model (SSM) architecture.
@@
 Multi-modal inputs (images via DINOv2, state) are fused and processed by sequential Mamba-2 blocks, updating the recurrent state $h_{t}$ which encodes history.
-At each step $t$ across the entire trajectory, MTIL predicts an action chunk $\hat{a}_{t:t+K-1}$ (current plus $K-1$ future steps).
+At each step $t$ across the entire trajectory, MTIL predicts an action chunk $\hat{a}_{t:t+K-1}$ (current plus K1K-1 future steps).
 This is supervised via L2 loss against ground truth actions $a_{t:t+K-1}$ from the demonstration (using last action for padding when near trajectory end).
 The historical context embedded in $h_{t}$ enables temporally coherent, long-horizon action generation.
-3.1 Background and Motivation
+III-A Background and Motivation
 Standard imitation learning often assumes a Markov Decision Process (MDP), learning reactive policies $\pi(a_{t}|o_{t})$ via Behavioral Cloning.
@@
 This motivates learning a compressed history representation $h_{t}\approx b_{t}$ using recurrent models.
-State Space Models (SSMs) like Mamba [16] offer a particularly compelling approach, providing a structured recurrent update $h_{t}=f(h_{t-1},x_{t})$ (where $x_{t}$ encodes $o_{t}$) with linear time complexity $O(L)$.
-This efficiency is critical for tractably encoding the long sequences required for full history representation, overcoming limitations of prior architectures [47].
+This aligns with the core ideas of both World Models, which learn a predictive latent state of the world, and Dynamical Systems (DS) approaches to control, which rely on an evolving internal state.
+State Space Models (SSMs) like Mamba [16] offer a particularly compelling synthesis of these ideas.
+They provide a structured recurrent update $h_{t}=f(h_{t-1},x_{t})$ (where $x_{t}$ encodes $o_{t}$) with linear time complexity $O(L)$.
+This efficiency is the critical enabler for tractably encoding the long sequences required for full history representation, overcoming the computational barriers of prior architectures [48].
 Our motivation for MTIL stems from leveraging Mamba’s state $h_{t}$ as this potent, efficiently computed representation of the full history.
+We view $h_{t}$ as the state of a learned, implicit dynamical system that acts as an action-oriented world model.
 By conditioning actions on both the current observation $o_{t}$ and this history-infused state $h_{t}$, MTIL learns a non-Markovian policy:
@@
 thereby directly addressing the core challenge of decision-making under ambiguity in POMDP-structured imitation learning by effectively utilizing the entire history.
-3.2 Leveraging Full Trajectory History with Mamba-2
+III-B Leveraging Full Trajectory History with Mamba-2
 MTIL employs Mamba-2 [49], an advanced structured State Space Model (SSM) notable for its refined selective mechanism and efficiency [16].
@@
 $y_{t}=\mathbf{C}_{t}h_{t}$
-Crucially, the input-dependent parameters $(\Delta_{t},\bar{\mathbf{B}}_{t},\mathbf{C}_{t})=f(x_{t})$ enable selective state dynamics, allowing the model to dynamically modulate information flow and preserve salient historical context.
-This selective mechanism, combined with inherent linear-time complexity $O(L)$, facilitates learning from complete trajectory histories—a significant advantage over quadratic-complexity $O(L^{2})$ attention mechanisms or memory-limited RNNs for long sequences.
+Crucially, the input-dependent parameters $(\Delta_{t},\bar{\mathbf{B}}_{t},\mathbf{C}_{t})=f(x_{t})$ enable selective state dynamics.
+This allows the model to learn a highly non-linear dynamical system where the system matrices themselves adapt based on the current input.
+This selective mechanism, combined with inherent linear-time complexity $O(L)$, facilitates learning from complete trajectory histories—a significant advantage over quadratic-complexity $O(L^{2})$ attention mechanisms.
 The resulting state $h_{t}$ acts as a dynamic summary of the salient history $(x_{1},...,x_{t-1})$, furnishing the requisite context even when the instantaneous observation $x_{t}$ is ambiguous.
@@
 Algorithm 1 MTIL Training (Sequential Step-based)
-0:  Expert trajectories $\mathcal{D}=\{\tau_{i}\}$, $\tau_{i}=(o_{1},a_{1},...,o_{T_{i}},a_{T_{i}})$, MTIL Policy $\pi_{\theta}$, initialized parameters $\theta$, Loss $\mathcal{L}$ (MSE), Chunk size $k$, Optimizer $Opt$
+0:  Expert trajectories $\mathcal{D}=\{\tau_{i}\}$, $\tau_{i}=(o_{1},a_{1},...,o_{T_{i}},a_{T_{i}})$, MTIL Policy $\pi_{\theta}$, initialized parameters $\theta$, Loss $\mathcal{L}$ (MSE), Chunk size $k$, Optimizer OptOpt
 1:  Initialize policy parameters $\theta$
@@
 10:  end for
-3.3 MTIL Training and Inference
+III-C MTIL Training and Inference
 MTIL enables imitation learning across complete expert trajectories, utilizing the architecture outlined in Figure 1.
 Distinctively, MTIL employs a sequential training procedure (Algorithm 1).
-This step-wise paradigm, leveraging Mamba’s recurrent ‘step‘ function, is fundamental to efficiently encoding arbitrarily long trajectory histories from high-dimensional observations (e.g., images) within feasible memory constraints—a key departure from parallel window-based approaches.At each timestep $t$ along the trajectory (length $T$), the policy receives the observation embedding $x_{t}$, updates its history-encoding hidden state from $h_{t-1}$ to $h_{t}$, and predicts a chunk of $K$ actions $\hat{a}_{t:t+K-1}$, covering the current and $K-1$ future steps.
-Learning proceeds by minimizing the Mean Squared Error (MSE) between this predicted chunk $\hat{a}_{t:t+K-1}$ and the ground truth actions $a_{t:t+K-1}$ sourced from the demonstration data.
-Notably, when $t+K-1$ exceeds the trajectory length $T$, the ground truth chunk is padded by repeating the final demonstrated action $a_{T}$.
-During inference (Algorithm 2), the trained policy operates autoregressively.
-It uses the same ‘step‘ function to update its hidden state $h_{t}$ based on the current observation and predict the action chunk $\hat{a}_{t:t+K-1}$.
-For enhanced stability and smoothness in execution, temporal aggregation strategies [19, 20, 5] are commonly applied.
-This involves computing the final action for time $t$ by applying a weighted average over multiple predictions made for this specific timestep but generated at different preceding steps.
-4 Experimental Results
-We conducted extensive experiments to evaluate the performance of MTIL across various benchmarks and real-world scenarios.See more details in Appendix A
-4.1 ACT benchmark
-Table 1: Success rates(%) on ACT.
-Cube Bimanual
-Method Transfer Insertion
-ACT [5] 90 50
-MTIL (10-step) 92 56
-MTIL (Full) 100 84
-MTIL’s efficiency and performance were evaluated on ACT [5] simulation tasks.
-Table 1 presents the final success rates, while Figure 2 shows learning curves and a backbone ablation.
-Compared to ACT [5], MTIL leveraging full trajectory history (approx.400 steps in these tasks) achieves substantially higher success rates (Table 1: 100%/84% vs.90%/50% for ACT on Cube Transfer/Bimanual Insertion) and markedly faster convergence (Figure 2a).
-We also evaluated MTIL variants using shorter fixed history windows (10, 20, and 50 steps),and their final performance did not show statistically significant differences from each other (10-step results: 92%/56%), confirming that such short-history modeling is clearly inferior to utilizing the full-history available.
-Furthermore, to isolate the benefits of the full-history approach from visual feature quality, we conducted an ablation comparing MTIL using DINOv2 [49] against ResNet18 [50] (which is used by ACT).
-While DINOv2 offers stronger features—a factor contributing to its selection, particularly for real-world robustness—the results (Figure 2b) combined with the short-versus-full history comparison underscore that the primary performance leap stems from MTIL’s core mechanism: encoding comprehensive trajectory context.
+This step-wise paradigm, leveraging Mamba’s recurrent ‘step‘ function, is fundamental to efficiently encoding arbitrarily long trajectories from high-dimensional observations (e.g., images) within feasible memory constraints—a key departure from parallel window-based approaches.A naive implementation of this sequential process would be limited to a batch size of one, posing a challenge for training efficiency.
+To address this, we introduce a novel batch-parallel training scheme.
+Instead of processing a single trajectory, our method processes a batch of $B$ trajectories simultaneously.
+At each timestep $t$, the model takes a batch of observations, updates their respective hidden states and computes the loss concurrently.
+This approach preserves the crucial temporal integrity within each trajectory while fully leveraging the parallel processing power of modern GPUs, making MTIL’s training time competitive with highly-parallelizable Markovian methods.During training, at each timestep $t$, the policy receives the observation embedding $x_{t}$, updates its history-encoding state from $h_{t-1}$ to $h_{t}$, and predicts an action chunk $\hat{a}_{t:t+K-1}$.
+Learning proceeds by minimizing the Mean Squared Error (MSE) against the ground truth actions $a_{t:t+K-1}$.
+During inference (Algorithm 2), the trained policy operates autoregressively, using the same ‘step‘ function to update its state and predict actions.
+For enhanced stability, temporal aggregation strategies [19, 20, 5] are applied, averaging over predictions from overlapping action chunks to produce a smoother final action.
+IV Experimental Results
+We conducted extensive experiments to evaluate the performance of MTIL across various benchmarks and real-world scenarios.All results stem from a rigorous protocol over three random seeds (100, 200, 300), with 50 roll-outs and the checkpoint for each run selected based on the minimum validation loss or as the final success rate for Robomimic.
+IV-A ACT benchmark
+TABLE I: Success Rates (%) on the ACT Benchmark. Results are averaged over 3 seeds, All experiments run on a single RTX 4090.
+Method History Length Cube Transfer (%) Bimanual Insertion (%)
+ACT [5] 1 (Markovian) 90.0 $\pm$ 2.0 50.0 $\pm$ 3.5
+Diffusion Policy [7] 1 (Markovian) 72.0 $\pm$ 2.6 28.0 $\pm$ 3.2
+Diffusion Policy [7] 10 78.0 $\pm$ 2.5 32.0 $\pm$ 4.1
+Diffusion Policy [7] 20 80.0 $\pm$ 2.2 34.0 $\pm$ 3.8
+Diffusion Policy [7] 30 82.0 $\pm$ 2.1 36.0 $\pm$ 3.5
+Diffusion Policy [7] 40 OOM
+Transformer-XL [32] Full (400) 86.0 $\pm$ 2.5 42.0 $\pm$ 4.0
+MTIL (10-step) 10 92.0 $\pm$ 1.5 56.0 $\pm$ 2.5
+MTIL (Full) Full (400) 100.0 $\pm$ 0.0 84.0 $\pm$ 2.1
 (a) Learning Curves
@@
 Figure 2: ACT benchmark performance.(a) Learning curves (Cube Transfer left, Bimanual Insertion right). (b) Backbone comparison (MTIL Full-History, DINOv2 vs. ResNet18).
-4.2 LIBERO Benchmark
+We evaluated MTIL on the ACT benchmark to dissect its performance, efficiency, and learning dynamics against prominent architectural paradigms.
+The results, which juxtapose success rates with architectural choices and history lengths, are presented in Table I.
+The findings decisively establish MTIL’s superiority.
+On both tasks, MTIL (Full) achieves a perfect or near-perfect success rate, drastically outperforming all baselines.
+The learning curves in Figure 2(a) illuminate this outcome, showing that MTIL not only attains a higher performance ceiling but also converges significantly faster, indicating a more stable and sample-efficient learning process.
+Conversely, the performance of attention-based models reveals a critical insight: naively incorporating history is an inefficient, and ultimately, a computationally infeasible strategy.
+While Diffusion Policy’s success rate scales with history length, it remains notably inferior to the simple Markovian ACT baseline and incurs a prohibitive computational cost, culminating in an Out-of-Memory (OOM) error.
+Even Transformer-XL, theoretically capable of full-history processing, fails to match ACT, reinforcing the hypothesis that attention is a suboptimal inductive bias for modeling the continuous dynamics of physical interaction.
+Furthermore, the backbone ablation in Figure 2(b) confirms our advantage is architectural.
+MTIL, even with an identical ResNet18 backbone [50], substantially outperforms ACT.
+The use of a stronger DINOv2 backbone [51] further widens this gap.This proves MTIL’s success stems from a fundamentally superior paradigm: a computationally efficient recurrent architecture that is intrinsically better suited to capturing the temporal fabric of the physical world.
+IV-B LIBERO Benchmark
 […]
-Table 2: Lifelong Learning Performance on LIBERO (EWC Strategy).
+TABLE II: Lifelong Learning Performance on LIBERO (EWC Strategy).
 LIBERO-LONG LIBERO-SPATIAL
@@
 ViT-M (Full) 0.58 $\pm$ 0.03 0.18 $\pm$ 0.04 0.25 $\pm$ 0.01 0.34 $\pm$ 0.04 0.10 $\pm$ 0.03 0.11 $\pm$ 0.01
-FWT($\uparrow$): Forward Transfer, NBT($\downarrow$): Backward Transfer, AUC($\uparrow$): Area Under Curve.
+FWT($\uparrow$): Forward Transfer, NBT($\downarrow$): Negative Backward Transfer (should be Backward Transfer, if it’s negative it’s good), AUC($\uparrow$): Area Under Curve.
 EWC strategy results averaged over 3 seeds (100, 200, 300) at 50 epochs.
@@
 Short-history (10-step,similar performance for 20/50 steps) and full-history result shown.
-4.3 Robomimic (Low-Dimensional State)
-To assess MTIL’s ability to model complex dynamics from proprioceptive states alone, we evaluated it on low-dimensional Robomimic tasks [12], comparing against prior methods including Diffusion Policy [7], Even without visual input, MTIL leveraging full history consistently surpasses previous state-of-the-art methods across the evaluated tasks.(Table 3).
-Table 3: Success Rates on Robomimic Low-Dimensional Tasks.
+IV-C Robomimic (Vision-based Policy)
+TABLE III: Behavior Cloning Benchmark (Visual Policy) on Robomimic. As per the original dataset, results are reported as final success rates.
 Lift Can Square Transport ToolHang
-Method ph mh ph mh ph mh ph mh ph mh
-LSTM-GMM [12] 1.00 0.93 1.00 0.81 0.95 0.59 0.76 0.20 0.67 0.31
-IBC [22] 0.79 0.15 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.00
-BET [26] 1.00 0.99 1.00 0.90 0.76 0.43 0.38 0.06 0.58 0.20
-DiffusionPolicy-C [7] 1.00 0.97 1.00 0.96 0.97 0.82 0.94 0.46 0.50 0.30
-DiffusionPolicy-T [7] 1.00 1.00 1.00 0.94 1.00 0.81 1.00 0.35 1.00 0.87
-MTIL (10-step) 1.00 0.94 1.00 0.83 0.96 0.61 0.78 0.22 0.70 0.33
-MTIL (Full History) 1.00 1.00 1.00 0.96 0.98 0.83 1.00 0.48 1.00 0.89
-An MTIL variant using only a 10-step history performs comparably, albeit slightly better, than LSTM-GMM [12], indicating limited gains from short recurrent contexts here.
-The full-history configuration, however, demonstrates the strength of Mamba’s architecture within MTIL for capturing intricate state transition dynamics and temporal dependencies inherent in manipulation, proving effective beyond visuomotor control.
-4.4 Real-World Dual-Arm Tasks
+ph mh ph mh ph mh ph mh ph
+LSTM-GMM [29] 1.00 1.00 1.00 0.98 0.82 0.64 0.88 0.44 0.68
+IBC [12] 0.94 0.39 0.08 0.00 0.03 0.00 0.00 0.00 0.00
+DiffusionPolicy-C 1.00 1.00 1.00 1.00 0.98 0.98 1.00 0.89 0.95
+DiffusionPolicy-T 1.00 1.00 1.00 1.00 1.00 0.94 0.98 0.73 0.76
+MTIL (10-step) 1.00 1.00 1.00 0.99 0.87 0.65 0.92 0.52 0.72
+MTIL (Full) 1.00 1.00 1.00 1.00 1.00 0.96 1.00 0.91 0.97
+To assess MTIL’s ability to handle high-dimensional visual inputs, we evaluated it on the vision-based Robomimic tasks [12].
+As shown in Table III, MTIL (Full) significantly outperforms all baselines, including the strong DiffusionPolicy variants.
+Notably, MTIL (10-step) offers only a marginal improvement over the LSTM-GMM baseline, highlighting that a short-history SSM is insufficient.
+The substantial performance gain of MTIL (Full) underscores its superior capability in leveraging full spatio-temporal context from visual data.
+This can be attributed to its nature as a learned dynamical system; the recurrent state $h_{t}$ acts as an implicit world model, tracking not just object locations but also their latent physical states (e.g., momentum, contact stability) over time, which is crucial for complex manipulation.
+IV-D Real-World Dual-Arm Tasks
 Figure 3: Dual UR3 experimental setup with four cameras (Top: Kinect; Side: D435i; Wrists: D405) and custom grippers.
+TABLE IV: Sequential Insertion Success Rates (%), averaged over 50 roll-outs.
+Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
+[…]
+[…]
+Figure 4: Sequential Insertion task stages (left panel of figure) and success rates (right panel, Table IV).
+MTIL successfully completes the sequence while ACT fails due to temporal ambiguity between stages 3 and 4.
+TABLE V: Coordinated Pouring Success Rates (%), averaged over 50 roll-outs.
+Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
+[…]
+[…]
+Figure 5: Coordinated Pouring task stages (left panel of figure) and success rates (right panel, Table V). MTIL achieves higher success and smoother execution compared to ACT.
 To validate MTIL in complex physical environments, we designed challenging tasks on a dual UR3 platform equipped with custom 2-finger grippers and four cameras providing multi-view observations (Figure 3).
 We compare MTIL (using DINOv2 backbone and full history) against ACT trained on identical demonstration data (100 demos per task).
+All real-world results are averaged over 50 evaluation roll-outs for the best checkpoint from each of the 3 seeds.
 Sequential Insertion Task.
-We designed this task (visualized in Figure 4) specifically to challenge Markovian policies by requiring long-term memory, a scenario where methods like ACT often fail.
+We designed this task (visualized in Figure 4) specifically to challenge Markovian policies by requiring long-term memory, a scenario where SOTA methods like ACT often fail.
 The four stages involve: (1) Left arm grasps Tube1, (2) Left passes Tube1 to Right arm, (3) Right arm inserts Tube1 into Tube2, (4) Right arm inserts Tube1 into Tube3.
 Critically, executing Stage 3 correctly necessitates recalling the completion of previous stages, as intermediate observations can be ambiguous.
-Table 4 details the stage-wise success rates.
+Table IV details the stage-wise success rates.
 MTIL, leveraging its full history state, successfully completes the entire sequence with high probability.
-In stark contrast, ACT, reliant on immediate context, is confounded by the temporal ambiguity; it frequently attempts Stage 4 directly after Stage 2, failing to execute the required sequence correctly and resulting in zero success for completing Stage 3, Stage 4, and the overall task.
+In stark contrast, ACT, reliant on immediate context, is confounded by the temporal ambiguity, as the observations after completing Stage 2 can be identical with completing Stage 3, making it indistinguishable for policies relying solely on current or short-term history.
+AS a result, it frequently attempts Stage 4 directly after Stage 2, failing to execute the required sequence correctly and resulting in zero success for completing Stage 3, Stage 4, and the overall task.
 This outcome underscores the limitations of short-history approaches and validates the imperative of encoding complete history for reliably executing temporally complex manipulation sequences.
-Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
-ACT 80 64 0 0 0
-MTIL 94 80 62 54 54
-Figure 4: Figure & Table 4: Sequential Insertion task stages (left) and success rates (%, right).
-MTIL successfully completes the sequence while ACT fails due to temporal ambiguity between stages 3 and 4.
-Method Stage 1 Stage 2 Stage 3 Stage 4 Overall
-ACT 80 64 50 32 32
-MTIL 94 80 74 62 62
-Figure 5: Figure & Table 5: Coordinated Pouring task stages (left) and success rates (%, right). MTIL achieves higher success and smoother execution compared to ACT.
 Coordinated Pouring Task.
@@
 While less susceptible to the specific ambiguity of the insertion task, it still requires accurate, temporally coordinated actions.
-Table 5 shows that although both methods achieve non-zero success, MTIL consistently outperforms ACT across the stages, resulting in a higher overall success rate and exhibiting notably smoother execution trajectories.
-5 Conclusion
+Table V (within Figure 5) shows that although both methods achieve non-zero success, MTIL consistently outperforms ACT across the stages, resulting in a higher overall success rate and exhibiting notably smoother execution trajectories.
+V Conclusion
 The trajectory of intelligence is intrinsically linked to the capacity for memory – the ability to weave the tapestry of past experiences into the fabric of present action.
-This work confronts a central limitation in contemporary imitation learning: the prevalent reliance on the Markovian assumption, which often reduces complex sequential behaviors to mere reactions to the immediate sensory world, neglecting the crucial dimension of history.
-We introduced Mamba Temporal Imitation Learning (MTIL), a paradigm shift that embraces the power of memory by leveraging the recurrent state dynamics inherent within the Mamba architecture.
-MTIL demonstrates that by encoding the full history of observations into a compressed, evolving state representation, robots can transcend the confines of the present moment.
-This comprehensive temporal context allows MTIL to effectively disambiguate perception and unlock the execution of intricate, state-dependent sequential tasks previously challenging for established methods like ACT and Diffusion Policy, as validated through extensive experiments across diverse simulated and real-world benchmarks.
-Our findings not only showcase the significant performance and efficiency gains afforded by MTIL but, more profoundly, underscore the essential role of history in bridging the gap between perception and intelligent, temporally coherent action in robotics.
-By demonstrating the efficacy of State Space Models in capturing the long flow of time, this work illuminates a promising pathway towards building robotic agents capable of deeper understanding and more sophisticated interaction with the world.
-Limitations
-While MTIL demonstrates significant advantages, we acknowledge certain limitations.
-Firstly, our current training strategy relies on sequential processing via Mamba’s ‘step‘ function(as detailed in Section 3.3).
-This is necessitated by the significant GPU memory demands of processing entire long sequences of high-dimensional observations concurrently, a challenge for current hardware.
-Unlike methods operating under the Markov assumption which might permit data shuffling, MTIL must preserve temporal order, inherently limiting parallelization compared to approaches like windowed Transformers and resulting in slower wall-clock training times per epoch—a trade-off for enabling full history modeling.
-Exploring techniques like gradient checkpointing or specialized hardware optimizations presents an avenue to mitigate this training speed limitation.
-Secondly, a potential failure mode, common to many imitation learning approaches, stems from distributional shift between training demonstrations and deployment scenarios.
-MTIL’s generalization capability to substantially novel environments or configurations may be limited, potentially leading to performance degradation when facing significant train-test mismatch.Lastly, while Mamba’s state aims to compress history effectively, its fidelity over ultra-long sequences (e.g., tens of thousands of steps) remains to be rigorously verified.
-Furthermore, performance can exhibit sensitivity to hyperparameter choices within the Mamba architecture and training regime, necessitating careful tuning.
-Addressing these aspects, alongside further theoretical investigation into the state representation limits, remains important future work.
+This work confronts a central limitation in contemporary imitation learning: the prevalent reliance on the Markovian assumption, which often reduces complex sequential behaviors to mere reactions to the immediate sensory world.
+We introduced Mamba Temporal Imitation Learning (MTIL), a new paradigm that embraces the power of memory by leveraging the recurrent state dynamics inherent within the Mamba architecture.
+We posit that MTIL represents a practical and powerful synthesis of concepts from World Models and Dynamical Systems.
+By encoding the full history of observations into a compressed, evolving state representation, MTIL learns an implicit, action-oriented world model.
+This comprehensive temporal context allows MTIL to effectively disambiguate perception and unlock the execution of intricate, state-dependent sequential tasks previously challenging for established methods.
+Our findings not only showcase the significant performance and efficiency gains afforded by MTIL but, more profoundly, underscore the essential role of history in bridging the gap between perception and intelligent action.
+By demonstrating the efficacy of SSMs in capturing the long flow of time in a computationally feasible manner, this work illuminates a promising pathway towards building robotic agents capable of deeper understanding and more sophisticated interaction with the world.
 Acknowledgments
-This work is supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.
+This work was supported by the Joint Funds of the National Natural Science Foundation of China (Grant No.
 U [^2] A20208), the Natural Science Foundation Innovation Group Project of Hubei Province (Grant No.
@@
 2022B0202010001-2).
-Appendix A Appendix
-A.1 ACT Benchmark Experimental Setup
-Tasks and Data.
-We evaluated MTIL on the two simulated tasks introduced in the ACT paper [5]: ”Cube Transfer” and ”Bimanual Insertion”.
-For both tasks, we used a dataset of 100 expert demonstration trajectories, automatically generated using scripted policies provided in the original ACT repository.
-The typical trajectory length for these tasks is around 400 steps.
-MTIL Configuration.
-The main MTIL model (”MTIL (Full)”) used full history encoding, processing the entire trajectory sequentially during training as described in Algorithm 1.
-We used an action chunk prediction size of $K=50$.
-The model was trained using the AdamW optimizer [53] with an initial learning rate of $2\times 10^{-4}$.
-A cosine learning rate decay schedule was applied over the training epochs.
-For the short-history ablation (”MTIL (10-step Hist.)”), the same architecture and training setup were used, but the hidden state was reset every 10 steps, limiting the effective history length.
-Backbones and Feature Processing.
-We experimented with two visual backbones to compare performance and isolate the effect of history modeling:
-  • ResNet18 (for fair comparison with ACT): Input RGB images (480x640x3) from multiple cameras were processed by separate ResNet18 [50] instances.
-Following the ACT pipeline, we extracted feature maps from an intermediate layer, resulting in 15x20x512 feature maps per camera.
-We applied learned positional encodings to these features.
-To enhance feature representation, we incorporated a Convolutional Block Attention Module (CBAM) [54] after the ResNet features, applying both channel and spatial attention mechanisms.
-The resulting features were flattened and projected using a linear layer to a common embedding dimension.
-  • DINOv2 (ViT-L/14, for best performance): Input RGB images (480x640x3) were fed into a pre-trained DINOv2 model with a ViT-L/14 architecture [49].
-The weights of the DINOv2 model were kept frozen during training.
-We extracted features from the output of the penultimate transformer block (specifically, ‘blocks[-4]‘ in common implementations), yielding a feature map of size approximately 34x45x1024 per camera (exact dimensions depend slightly on implementation details of patch embedding and image size).
-These high-dimensional features were then flattened and projected via a linear layer to the common embedding dimension.
-Multi-Modal Fusion.
-The projected visual features from all cameras were concatenated.
-Low-dimensional robot state information (e.g., joint positions, gripper width/status) was also concatenated into a vector.
-To effectively integrate these modalities, we employed a cross-modal attention mechanism: the concatenated visual features served as query, while the low-dimensional state vector served as the keys and values.
-The output of this attention mechanism was then passed through a final linear layer to project it to the input dimension ($D$) required by the Mamba-2 blocks.
-Mamba-2 Architecture.
-The core sequence modeling was performed by a stack of Mamba-2 blocks.
-Key hyperparameters such as the model dimension ($d_{model}$), state expansion factor, state dimension ($d_{state}$), and the number of layers were tuned based on validation performance.
-Specific values are provided in the codebase and Table 6.
-A.2 LIBERO Benchmark Experimental Setup
-Goal and Setup.
-The experiments on LIBERO [51] aimed to evaluate MTIL’s performance in a challenging lifelong learning setting, focusing on knowledge transfer and catastrophic forgetting.
-To ensure fair comparison with the baselines reported in [51], we adhered closely to their experimental protocol.
-Backbones and Prediction.
-We used the standard ResNet and ViT [55] backbones provided within the official LIBERO benchmark codebase.
-We did *not* use the DINOv2 backbone in these experiments to maintain comparability.
-Furthermore, following the LIBERO setup for RNN/Transformer baselines, MTIL was configured to predict only the action for the next single step ($K=1$), rather than action chunks.
-Language Encoder and Fusion.
-Task descriptions were encoded using the standard BERT [56] encoder, consistent with the LIBERO baselines.
-For fusion, simple concatenation of visual features, state information, and language embeddings followed by linear projection was used, matching the baseline setup.
-No cross-modal attention mechanism was employed here.
-Lifelong Learning Strategy: EWC.
-We employed the Elastic Weight Consolidation (EWC) [52] strategy to mitigate catastrophic forgetting as tasks were learned sequentially.
-EWC works by calculating the Fisher Information Matrix (FIM) for the parameters of the model with respect to previously learned tasks.
-During training on a new task, a quadratic penalty term is added to the loss function, discouraging changes to parameters deemed important for past tasks (i.e., those with high Fisher information).
-Rationale for EWC with MTIL: We selected EWC not only because it is a standard lifelong learning technique used in the benchmark, but also due to its theoretical synergy with MTIL’s state-based history encoding.
-The Mamba-2 blocks contain parameters that govern the selective state updates ($\Delta_{t},\mathbf{B}_{t},\mathbf{C}_{t}$) and state transitions ($\mathbf{A}_{t}$).
-These parameters are crucial for capturing the temporal dynamics and historical context learned from demonstrations.
-By applying EWC, we hypothesize that the FIM helps identify and protect those parameters within the Mamba architecture that are most critical for encoding the temporal patterns learned in previous tasks.
-This allows MTIL to retain its ability to leverage history effectively for older tasks (leading to low NBT) while still adapting its state dynamics to accommodate new tasks (enabling high FWT).
-Metrics Definition.
-We report the standard LIBERO lifelong learning metrics, averaged over 3 random seeds after training on all tasks in a suite for 50 epochs using the EWC strategy:
-  • FWT (Forward Transfer, $\uparrow$): Measures the influence of having learned tasks $11,...,i-$ 1 on the initial performance of learning task $i$, compared to learning task $i$ from scratch.
-It is calculated as:
-    $FWT=\frac{1}{N-1}\sum_{i=2}^{N}(A_{i-1,i}-A_{base,i})$
-    where $A_{i-1,i}$ is the accuracy on task $i$ evaluated *immediately after* training on task $i-1$ (before training on task $i$), and $A_{base,i}$ is the accuracy on task $i$ when trained in isolation (from scratch).
-Higher values indicate better positive transfer.
-(Note: LIBERO paper definition might vary slightly, we follow the standard interpretation of immediate transfer).
-  • NBT (Negative Backward Transfer, $\downarrow$): Measures the average drop in performance on previously learned tasks after learning subsequent tasks (i.e., forgetting). It is calculated as:
-    $NBT=\frac{1}{N-1}\sum_{i=1}^{N-1}(\max(0,A_{i,i}-A_{N,i}))$
-    where $A_{i,i}$ is the final accuracy on task $i$ after training on task $i$, and $A_{N,i}$ is the accuracy on task $i$ after training on all $N$ tasks. Lower values indicate less forgetting.
-  • AUC (Area Under the learning Curve, $\uparrow$): Represents the overall performance across all tasks throughout the lifelong learning process.
-It is typically calculated as the average accuracy across all tasks, evaluated after the model has finished training on the final task $N$:
-    $AUC=\frac{1}{N}\sum_{i=1}^{N}A_{N,i}$
-    Higher values indicate better overall performance retention and learning.
-A.3 Robomimic Benchmark Experimental Setup
-Goal and Setup.
-The experiments on Robomimic [12] aimed to evaluate MTIL’s ability to model complex manipulation skills using only low-dimensional proprioceptive state information (e.g., joint positions, velocities, end-effector pose, gripper status), without any visual input.
-Input and Settings.
-We used the standard low-dimensional state datasets provided by Robomimic for tasks like Lift, Can, Square, Transport, and Tool Hang, considering both proficient-human (ph) and multi-human (mh) data sources.
-To ensure fair comparison, we followed the evaluation protocol and data splits used in prior work, particularly matching the setup for the LSTM-GMM baseline [12] where applicable.
-Key Difference and Action Output.
-The primary difference between MTIL and the LSTM-GMM baseline lies in the history modeling.
-While LSTM-GMM uses LSTMs, often trained on shorter sequences (e.g., 10-50 steps) in practice [12], MTIL (Full History) processes the entire demonstration sequence using its Mamba-based recurrent state.
-To isolate the effect of the sequence modeling backbone, we configured MTIL to use a Gaussian Mixture Model (GMM) head for predicting the action distribution at each step, mirroring the output structure of LSTM-GMM.
-The number of mixture components was treated as a hyperparameter.
-MTIL Architecture.
-A stack of Mamba-2 blocks was used to process the sequence of low-dimensional state vectors.
-The specific hyperparameters (e.g., $d_{model}$, $d_{state}$, number of layers) were optimized for these tasks (See details in Table 6).
-A.4 Real-World Experimental Setup
-Figure 6: Details of the experimental hardware, showing the dual-arm robot setup (UR3) with the task workspace, and the specific custom gripper, wrist camera(D405), Kinect camera and RealSense D435 utilized.
-Platform.
-Experiments were conducted on a physical setup consisting of two Universal Robots UR3 arms mounted on a table.
-Each arm was equipped with a custom 3D-printed 2-finger parallel gripper actuated by a Dynamixel servo motor.
-The sensory input comprised multi-view images from four cameras: one Intel RealSense D435i providing a side view, one Microsoft Azure Kinect providing a top-down view, and two Intel RealSense D405 cameras mounted on the wrists of the robots for close-up views (see Figure  6).
-MTIL Configuration.
-For the real-world tasks (Sequential Insertion, Coordinated Pouring), we used the full MTIL configuration designed for robust performance: full history encoding via sequential training (Algorithm 1), the DINOv2 (ViT-L/14) backbone with frozen weights for visual feature extraction, and prediction of action chunks with $K=50$.
-Multi-modal fusion using cross-attention between visual features and low-dimensional state (joint positions, gripper state) was employed, identical to the setup described in Appendix A.1.
-Specific Mamba-2 architecture parameters were consistent with those used in the ACT benchmark simulation experiments with DINOv2.
-Data Collection.
-For each of the two tasks, we collected 100 expert demonstrations using a custom teleoperation system that use spacemouse.
-The system allowed a human operator to control the dual arms simultaneously via input devices, recording the robot states and camera images at each timestep.The trajectory length for the Sequential Insertion task is around 500 steps while the trajectory length for the Coordinated Pouring task is around 1000 steps.
-Inference Details: Chunking and Temporal Aggregation.
-During policy execution on the real robot (inference time, Algorithm 2), we applied both action chunking (predicting $K=50$ steps ahead) and temporal aggregation.
-Temporal aggregation was implemented using a weighted averaging scheme over overlapping predicted chunks.
-Specifically, the action $a_{t}^{\text{final}}$ executed at time $t$ was computed as a weighted average of predictions made for time $t$ from chunks initiated at previous times $j$ (where $j\leq t<j+K$).
-We used exponentially decaying weights $W=\{w_{0},...,w_{K-1}\}$ where $w_{k}\propto\gamma^{k}$ for some decay factor $\gamma<1$, giving higher importance to predictions made closer to the current time step (i.e., smaller $k=t-j$).
-Rationale: Applying chunking and temporal aggregation is a common practice in methods like ACT [5] and has been shown to significantly improve the smoothness and stability of executed robot motions by reducing jitter [19, 20].
-We employed this technique primarily for two reasons: (1) To ensure a fair comparison with ACT, which also utilizes this strategy.
-(2) To achieve better practical performance and robustness during real-world execution, as smoother actions are generally preferable.
-While MTIL’s inherent temporal modeling capability provides a strong foundation for coherent action sequences, the temporal aggregation acts as an additional filter that further enhances stability, leading to more reliable task execution.
-A.5 Hyperparameter Summary
-Table 6: Key MTIL Hyperparameters Across Benchmarks.
-Parameter ACT Sim LIBERO Robomimic Real World
-Backbone ResNet18 / DINOv2 ResNet / ViT N/A (State) DINOv2
-History Full / 10-step Full / 10-step Full / 10-step Full
-Chunk Size (K) 50 1 1 50
-Mamba $d_{model}$ 2048 256 64 2048
-Mamba $d_{state}$ 512 64 16 512
-Mamba Layers 4 1 4 4
-Optimizer AdamW AdamW AdamW AdamW
-Initial LR 2e-4 1e-4 1e-4 2e-4
-LR Schedule Cosine Cosine Cosine Cosine
-Cross-Attention Yes No N/A Yes
-Output Head Linear Linear GMM Linear
-Temporal Aggregation Yes No No Yes